OpenStudy (anonymous):

can someone explain how to find two positive numbers such that their product is 192 and the sum of the first plus three times the second is a minimum?

hartnn (hartnn):

let 2 numbers be x and y 'their product is 192 ' will give you one equation, can u find that ?

OpenStudy (anonymous):

x+3y=192?

hartnn (hartnn):

product means xy

OpenStudy (anonymous):

i can get it to where it equals 192y-3y^2 but i don't know where to go from there to find the two numbers..

hartnn (hartnn):

xy=192---> x=192/y f(y) = x+3y = 192/y +3y now differentiate f(y) and equate it to 0, and tell me what u get.